Bronson Paper Products purchased 17,500 acres of forested timberland in March 2014. The company paid $2,975 per acre for this land, which was above the $1,400 per acre most farmers were paying for cleared land. During April, May, June, and July 2014, Bronson cut enough timber to build roads using moveable equipment purchased on April 1, 2014. The cost of the roads was $437,500, and the cost of the equipment was $393,750; this equipment was expected to have a $15,750 salvage value and would be used for the next 15 years. Bronson selected the straight-line method of depreciation for the moveable equipment. Bronson began actively harvesting timber in August and by December had harvested and sold 945,000 board feet of timber of the estimated 11,812,500 board feet available for cutting.

In March 2015, Bronson planted new seedlings in the area harvested during the winter. Cost of planting these seedlings was $210,000. In addition, Bronson spent $14,000 in road maintenance and $10,500 for pest spraying during calendar-year 2015. The road maintenance and spraying are annual costs. During 2015, Bronson harvested and sold 1,354,500 board feet of timber of the estimated 11,287,500 board feet available for cutting.

In March 2016, Bronson again planted new seedlings at a cost of $262,500, and also spent $26,250 on road maintenance and pest spraying. During 2016, the company harvested and sold 1,137,500 board feet of timber of the estimated 11,375,000 board feet available for cutting.

Compute the amount of depreciation and depletion expense for each of the 3 years (2014, 2015, 2016). Assume that the roads are usable only for logging and therefore are included in the depletion base.
Answer #1

Year 2014:

Depletion base – Timber

Cost of timber = 2975-1400(land cost)*17500 = 2,75,62,500

Cost of road    437,500

               Depletion base 2,80,00,000

Depletion to be charged for the year = (28000000/11812500)*945000= $2240000

Removable Equipment:

Depreciation (SLM) = (393,750-15750)/15 = $25200.

Proportionate depreciation =25200*9/12 = $18900.

Year 2015:

Timber depletion base:

Balance after depletion in 2014    =    25760000

Add: cost of planting seeds 210000

Depletion base for the year 25970000

Depletion = (25970000/11,287,500)*1,354,500 = $31,16,400

Depreciation for the removable equipment   = $25,200.

Year 2016:

Timber:

Depletion base b/f                = $2,28,53,600

Add cost of planting seeds =    262,500

Depletion base for the year=    2,31,16,100

Depletion for the year = (23116100/11,375,000)*1,137,500 = $23,11,610.

Depreciation on the removable equipment                   = $25,200.